OpenStudy (anonymous):

why does this system cannot be solved? 2x + 2y + 3z = 23 2x + 2y + 5z = 33 -4x - 4y - 8z = -55

OpenStudy (anonymous):

Subtract first equation from second equation gives 2z = 10 or z = 5. If you multiply the first equation by 2 and add it to the third equation, you get -2z = 9 or z = -9/2. We have two different values for z, so the system can not have a solution.

OpenStudy (anonymous):

No, I mean, there is a way that allows me at first glance to figure out that this equation has not a solution?

OpenStudy (anonymous):

Add the first 2 equations and multiply the result by -1. You get -4x -4y -8z = -56, but the third equation is -4x -4y -8z = -55

OpenStudy (anonymous):

lol, I misunderstood a question before so I thought there was a simple way to see if one system has/has not a solution. I know how to solve that system. Thank you very much djvu